OpenStudy (anonymous):

The equation of a circle is (x + 3)2 + (y + 5)2 = 8. Determine the coordinates of the center of the circle.

OpenStudy (anonymous):

The equation of a circle with center (h, k) and radius "r" is:\[(x - h)^{2} + (y - k)^{2} = r ^{2}\]Match up the "h" and the "k" from the equation to the representation of the "h" and the "k" in the ordered pair and be careful about the negative signs! That means that if you don't have a negative number for both "h" and "k", you have a wrong answer right off the bat.

OpenStudy (anonymous):

Wait so do I put +or - sorry I'm comfused

OpenStudy (anonymous):

since -h is the x cordinate h = -3 so h= 3 and k = -5

OpenStudy (anonymous):

If you are not sure about the sign, get your "h" and your "k" and then put them into the general equation and see if the equation resolves down to the equation you are given. If the signs don't look right, just repeat the process with a different sign.

OpenStudy (anonymous):

(X-3)2+(y-5)2=8^2

OpenStudy (anonymous):

Try h = -3 and k = -5 That is (-3, -5) What do you get when you put that into the general equation? Does it match your original equation? Remember, when you subtract a negative, it is the same as adding a positive: x - (-3) = x + 3

OpenStudy (anonymous):

(X+3)2+(y+5)=8^2

OpenStudy (anonymous):

Close. The center is (-3, -5) but the right side is [sqrt(8)]^2 = 8 not 8^2

OpenStudy (anonymous):

Ohh ok

OpenStudy (anonymous):

Anyway the question was only asking for the center, and now you have that for your answer.

OpenStudy (anonymous):

All good now, @nikakuz ?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @nikakuz